OpenStudy (anonymous):

Solve by elimination 1. 2x-3y=61 2x+y=-7 2. 2x+18y=-9 4x+18y= =27 3. 3x-10y=-25 4x+40y= 20 4. x-8y=18 -16x+16y=-8 5. 88x-5y=39 -8x+3y=-1

OpenStudy (anonymous):

for first one x=12 and y=-17

OpenStudy (anonymous):

Explain please.

OpenStudy (anonymous):

just sbtract first equation from second one those equations u can eliminate one variable so that u can find one by substituion u can also find the other one u can do the other problems in the same way! if u want any help reply!

OpenStudy (anonymous):

i m doing the problem no.1 and u will do the rest problems in a similar manner it is given 2x+y =-7 therefore y = -7-2x now plug in the above value of y in 2x-3y = 61 2x-3(-7-2x) = 61 2x+21+6x = 61 8x= 61-21 8x = 40 x = 5 now plugin the value of x=5 in 2x+y = -7 2(5)+y = -7 10+y = -7 y = -10-7 y =-17

OpenStudy (anonymous):

We just learned that way. Now we're learning how to do it with elimination so that's not the way she wants it done.